How much water must be evaporated from 200 mL of a 5% salt solution to produce a 25% salt solution?
Answer by stanbon(75887)

How much water must be evaporated from 200 mL of a 5% salt solution to produce a 25% salt solution?
----
Equation:
water - water = water
0.95*200 - 1*x = 0.75(200-x)
----------------------------
95*200 - 100x = 75*200 - 75x
-----
25x = 20*200
x = (4/5)200 = 160 mL (amt. of water to evaporate)
----------
Cheers,
Stan H.
------------